Given three consecutive integers. When half the secondinteger is subtracted from twice the first one the result is eleven less than the third one. Find the three integers.
Answer by checkley77(12569)

Let x, x+1 & x+2 be the 3 integers.
2x-(x+1)/2=(x+2)-11
2x-x/2-1/2=x+2-11
3x/2-x=-9+1/2
x/2=-17/2
x=-17 Answer for the smallest integer.
-17+1=-16 for the middle integer.
-17+2=-15 for the larger integer..
Proof:
2*-17-(-16/2)=-15-11
-34-(-8)=-26
-34+6=-26
-26=-26